OpenStudy (anonymous):

limit as x->-∞ from (4x^3+2)/(x^3+√(4x^6+1) I know is -4 I'd like to see how you guys do it .

OpenStudy (zarkon):

factor out an \(x^3\) from the top and bottom.

OpenStudy (zarkon):

notice \[\sqrt{x^6}=|x^3|\]

OpenStudy (turingtest):

\[\lim_{x\to-\infty}{4x^3+2\over x^3+\sqrt{4x^6+1}}\]yup easiest way^

OpenStudy (anonymous):

How do you isolate \(x^3\) from \(\sqrt{4x^6+1}\). All I see is\[\sqrt{x^6(4+x^{-6})}\]

OpenStudy (anonymous):

I would've suggested multiplying by the reciprocal.

OpenStudy (zarkon):

@badreferences use my second post

OpenStudy (anonymous):

I still don't see it.\[\sqrt{4x^6+1}=\sqrt{x^6\left(4+x^{-6}\right)}=\left|x^3\right|\sqrt{4+x^{-6}}\]If you don't mind me asking.

OpenStudy (zarkon):

right...and if \(x<0\) then \(|x^3|=-x^3\)

OpenStudy (anonymous):

4/(1-2) ..

OpenStudy (anonymous):

oh, okay, I see it. then afterwards multiply by \[1+\sqrt{4+x^{-6}}\] reciprocal

OpenStudy (zarkon):

you don't need to

OpenStudy (anonymous):

oh, find the limit. not simplify. sorry, wasn't reading.

OpenStudy (zarkon):

just let \(x\to-\infty\)

OpenStudy (anonymous):

my problem ... |x^3| = -x^3

OpenStudy (zarkon):

\[|x|=\left\{\begin{matrix}x & if & x\ge 0 \\ -x & if & x<0\end{matrix}\right.\]

OpenStudy (zarkon):

replace \(x\) by \(x^3\)

OpenStudy (anonymous):

right ...

OpenStudy (anonymous):

THX

OpenStudy (zarkon):

no problem